# Robust Pronoun Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

### v2

@@ Abstract
Robust, faithful and harm-free pronoun use for individuals is an important goal for language models as their use increases, but prior work tends to study only one or two of these characteristics at a time. To measure progress towards the combined goal, we introduce the task of pronoun fidelity: given a context introducing a co-referring entity and pronoun, the task is to reuse the correct pronoun later. We present RUFF, a carefully-designed dataset of over 5 million instances to measure robust pronoun fidelity in English, and we evaluate 37 popular large language models across architectures (encoder-only, decoder-only and encoder-decoder) and scales (11M-70B parameters). When an individual is introduced with a pronoun, models can mostly faithfully reuse this pronoun in the next sentence, but they are significantly worse with she/her/her, singular they and neopronouns. Moreover, models are easily distracted by non-adversarial sentences discussing other people; even one additional sentence with a distractor pronoun causes accuracy to drop on average by 34%. Our results show that pronoun fidelity is neither robust, nor due to reasoning, in a simple, naturalistic setting where humans achieve nearly 100% accuracy. We encourage researchers to bridge the gaps we find and to carefully evaluate reasoning in settings where superficial repetition might inflate perceptions of model performance.
@@ 1 Introduction
Accordingly, it is important for large language models (LLMs) to use pronouns faithfully and without causing harm. To this end, many studies have explored how LLMs handle pronouns, showing that they stereotypically associate pronouns and occupations (Kurita et al., 2019), reason about co-referring pronouns and entities better when they conform to stereotypes (Tal et al., 2022), fail when exposed to novel pronoun phenomena such as neopronouns (Lauscher et al., 2023), and cannot consistently reuse neopronouns during generation (Ovalle et al., 2023a). These shortcomings create quality of service differences and cause representational harm, amplifying discrimination against certain pronoun users (Blodgett et al., 2020; Dev et al., 2021).
However, a question that has gone unexamined thus far is: How robust is model faithfulness to pronouns when discussing more than one person? To answer this question, we propose pronoun fidelity2), a new task to investigate realistic model reasoning about pronouns, and we introduce RUFF3), a novel, large-scale dataset of over 5 million instances, carefully designed to evaluate this task. With this dataset, we present an analysis of pronoun fidelity across 37 popular language models covering a range of architectures and scales (11M-70B), to investigate whether models are reasoning, repeating, or just biased.
@@ 3 RUFF Dataset
RUFF, our evaluation dataset for Robust pronoUn Fidelity at scale, is constructed with a 3-step pipeline to create naturalistic instances of our task with unambiguous answers: template creation (§3.1), assembly (§3.2), and validation (§3.3).
Our data covers 60 occupations (see Appendix B) and four third-person pronouns (he, she, they and xe) in three grammatical cases (nominative, accusative and possessive dependent). Our occupations (and corresponding participants for distractor sentences) are chosen from Winogender schemas (Rudinger et al., 2018), as their bias characteristics are well-studied in NLP. In addition to the English masculine (he/him/his) and feminine (she/her/her) pronouns, we heed Lauscher et al.’s (2022) call for more inclusive NLP research by examining two more pronoun sets that are less well-studied in NLP: singular they (they/them/their), the pronoun of choice of over 75% of respondents to the Gender Census (Lodge, 2023), and xe/xem/xyr, the most popular neopronoun according to the same census.
@@ 3.1 Template creation
Task templates. We create one task sentence template per occupation and grammatical case for a total of 180 templates, designed to create an unambiguous coreference with the occupation only. For instance, charges for preparing tax returns can only belong to an accountant, never a taxpayer, which is the corresponding participant.
Context templates. We create explicit (definite reference + pronoun) and implicit (pronoun-only) versions of 10 context templates per grammatical case, for a total of 30 templates. Each explicit context template begins with an entity and introduces the pronoun in a clause, e.g., The taxpayer needed coffee because their day had started very early, while implicit templates are simple sentences like Their sleep had been fitful. All templates contain generic themes, e.g., universal human emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.). These templates can thus be applied to all occupations and participants, and allow for a controlled but still coherent setting.
@@ 3.2 Template assembly
We then instantiate and combine the previously created templates to assemble our data instances. First, we select an occupation ($e_{a}$) and one of its task templates. We pick a pronoun ($p_{a}$) to use as ground truth and instantiate a random context template with the selected occupation and pronoun. The simplest version of the pronoun fidelity task includes just this introduction sentence followed by the task sentence. Instantiating 10 templates with 4 different pronoun sets and pairing them with task templates for 60 occupations across 3 grammatical cases gives us a total of 7,200 unique instances for this version of the task.
To create more complex data instances, we insert a variable number of distractor sentences between the introduction and task sentences, discussing a participant $e_{b}$ who uses a different pronoun $p_{b}$. These are also sampled from the set of context templates (see Appendix C for details). Instantiating 4 templates with 3 unused pronouns gives 86,400 unique instances with one distractor.
@@
FLAN-T5 small (77M), base (248M), large (783), xl (2.85B), xxl (11.3B) Encoder-decoder
@@
Table 2: Models we experiment with across a range of sizes and architectures.
@@ 3.3 Data validation
We validate all task and context templates. To verify that the pronoun fidelity task is easy and unambiguous for humans, we also validate a subset of these instances with 0-5 distractors. Annotator information is shown in Appendix D and all annotator instructions are provided in Appendix E.
@@ 4 Experimental Setup
We list our models, evaluation methods, and metrics. Further details are provided in Appendix F.
@@ 4.1 Models
We experiment with 37 transformer-based language models (see Table 2), chosen to evaluate the effects of architecture and scaling. Our 9 encoder-only models are from the BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), ALBERT-v2 (Lan et al., 2020) and MosaicBERT (Portes et al., 2023) model families, as the first three remain well-used in NLP, and the last is trained on much more data. As for our 20 decoder-only models, we select the popular Llama-2 (Touvron et al., 2023) model family, as well as OPT (Zhang et al., 2022) and Pythia (Biderman et al., 2023) for their large range of model sizes. In Appendix I, we also experiment with eight popular chat models that are further trained with instruction-tuning and reinforcement learning, to evaluate task performance with prompting; specifically, we use decoder-only Llama-2-chat models (Touvron et al., 2023) and encoder-decoder FLAN-T5 models (Chung et al., 2022).
@@ 4.2 Obtaining predictions
Decoder-only models. For the majority of our experiments, we follow Hu and Levy (2023) in taking direct measurements of probabilities as a proxy for models’ metalinguistic judgements. We verbalize four versions of each data instance, i.e., we fill in the blank with each of the four pronouns we consider, creating four options. We then compute instance-level model log likelihoods for these options, and select the option with the highest log likelihood as the model’s choice.
Encoder-only models. For comparability with decoder-only models, we follow the same procedure for encoder-only models, but use pseudo log likelihoods (Salazar et al., 2020; Kauf and Ivanova, 2023). We do not use masked token prediction due to tokenization issues with neopronouns (Ovalle et al., 2023b); briefly, we want xe to be tokenized “normally” (which is often as two tokens) rather than a single UNK token.
Chat models. Following common practice, we evaluate chat models (FLAN-T5 and Llama-2-chat) using vanilla and chain-of-thought prompting in Appendix I. Following Sclar et al. (2024), we show the range of expected performance with 10 different prompts (see Appendix G), inspired by the prompts to elicit coreferences in the FLAN collection (Longpre et al., 2023).
@@ 5 Model Predictions with No Context
As we cannot evaluate accuracy on a task with no single correct answer, we show the counts of model predictions of different pronoun declensions in Figure 2, averaged over all models. Model-specific counts are shown in Appendix H. Even though our task sentences are designed such that any pronoun set can be used grammatically, all models tend to assign higher probability to he/him/his than other pronoun sets.
@@ 6 Injecting an Introductory Context
Which pronouns are harder? Even in the simplest case of the pronoun fidelity task, patterns emerge when split by pronoun, as shown in Figure 4. Overall model accuracy on he/him/his is significantly higher than she/her/her, which in turn is significantly higher than both they/them/their and xe/xem/xyr, in line with previous findings that language technology has gaps when it comes to neopronouns (Lauscher et al., 2023). Models show intriguing patterns with these last two pronoun sets. Most encoder-only models appear to handle the neopronoun better than singular they (e.g., BERT-large has an accuracy of 0.78 on xe/xem/xyr compared to 0.60 on they/them/their), which warrants further investigation. Decoder-only models smaller than 6.7B parameters struggle with the neopronoun, with every OPT and Pythia model smaller than 2.7B parameters performing below chance, and in some cases (e.g., Pythia-14M, Pythia-70M and Pythia-160M) even performing close to 0.0. Beyond this scale, however, models perform better on xe/xem/xyr than on singular they, with Llama-13B achieving 0.96 accuracy on the neopronoun. These differences are statistically significant. As the training data for individual model families is the same, this might suggest that decoder-only models generalize to novel pronouns starting at the scale of 6.7B parameters, but in light of Schaeffer et al. (2023), this result could just as well be a mirage resulting from our use of accuracy, a discontinuous metric. In either case, our observations could also explain the poor performance that some previous studies of neopronouns find, as the largest model that Hossain et al. (2023) experiment with, for instance, is OPT-6.7B. The lower performance of bigger models with singular they could also be a reflection of human processing difficulties with definite, specific (singular) they, as has been observed in linguistics (Conrod, 2019).
@@ 7 Adding Distractors
Using vanilla and chain-of-thought prompting (Appendix I) show the same patterns of degradation, reinforcing that model pronoun fidelity is not robust, and good performance with no distractors (§6) is likely not due to “reasoning” at all.
@@ 8 Distractibility versus Bias
In adding distractor sentences, we add distance from the introduction via additional tokens that might make the model forget the original occupation-pronoun association, and the distractor pronoun also acts as a competing token that the model might accidentally repeat. In this section, we focus on the error cases to disentangle whether models are “forgetting” and reverting to biased predictions from Section 5, or if they are actually being distracted. When a model gets the answer wrong, it is for one of three reasons: (1) distractibility, i.e., repeating the distractor pronoun, (2) bias, i.e., reverting to the model’s context-free prediction, or (3) picking a completely different pronoun. All three possibilities are illustrated below , and we hypothesize that the first two possibilities are much more frequent than the third.
@@
We first examine model distractibility, i.e., what percentage of errors are caused by models repeating the distractor pronoun instead of the correct pronoun. As expected, Figure 7 shows that across models, distraction is indeed the primary type of error for most models. Decoder-only models get increasingly distractible with more distractors, i.e., the proportion of errors due to distractor pronoun repetition steadily increases as distractors are added, saturating just below 85%. On the other hand, encoder-only models seem to become less distractible with the addition of more distractors. We know from the previous section that encoder-only models recover in their pronoun fidelity with 2-5 distractors, but here we measure distractibility as a percentage of all errors. Thus, a constant or increasing proportion of all the model errors could be due to distraction, and the fact that it isn’t for encoder-only models is quite surprising! We leave it to future work to investigate whether this behaviour relates to positional bias or context use.
@@
The high distractibility of all models strongly suggests that models simply repeat the most recent pronoun, rather than using anything resembling “reasoning,” and the contrast in error behaviour between encoder-only and decoder-only models further highlights their differences. This shows that claims about decoder-only models should not be applied to all LLMs, and that reasoning must be evaluated carefully, accounting for the possibility of inflated performance due to shallow heuristics like repetition.
@@ 9 Discussion and Future Work
Our results show that even the biggest models of today are not up to the task of pronoun fidelity once it includes a single sentence discussing another person. All models are easily distracted, but encoder-only models and decoder-only models show very different patterns both in performance degradation with more distractors and their reasons for errors. Performance on this type of reasoning task should be evaluated carefully, with attention to how the overall patterns break down by different pronouns and accounting for the possibility of repetition. Below we expand on some questions raised by our findings.
@@
Improving robust pronoun fidelity. A natural direction of future work is to solve the problem of robust pronoun fidelity. We urge researchers interested in this direction to treat RUFF as an evaluation dataset, as it was designed. Due to the presence of positional and associative heuristics that we expand on in Appendix A, RUFF should not be seen as a source of data for fine-tuning or in-context learning, which is also why we do not run these experiments. Instead, a promising direction might be to encourage models to explicitly track associations between pronoun sets and individuals, just as people do.
@@
On “reasoning.” Throughout the paper, we refer to “reasoning,” but this is inaccurate. What looks like “reasoning” when we inject an introductory context starts to look much more like repetition—or stochastic parroting (Bender et al., 2021)—when we add a distractor and see the same models performing drastically worse. Even the higher performance of encoder-only models cannot accurately be attributed to “reasoning” in the same way that we use this word for humans, as these models are not grounded in meaning from the real world (Bender and Koller, 2020). We use the word reasoning in line with other work in the field, but note that as these are all language models, it is more accurate to say that the way that decoder-only models model language is prone to repetition of recent examples of the same word class, compared to encoder-only models.
@@
Why exactly do we see the patterns we see? Our dataset design and error analysis shed light on model behaviour, allowing us to evaluate different architectures comparably and disentangle the effects of repetition, distraction and statistical bias. However, it is beyond the scope of this paper to investigate where in the model architecture, neurons or pre-training data this comes from and what we can do about it towards improving reasoning and mitigating bias. Tools from interpretability literature, e.g., attribution analysis, could help here, and are an important direction for future work.
@@
Beyond our dataset. Given the breadth of our task definition, future work could include examining pronoun fidelity for participants, for names by extending Hossain et al. (2023), with differently ordered sentences, and in more natural settings such as stories (Mostafazadeh et al., 2016) or with real-world data as in Webster et al. (2018) and Levy et al. (2021). Additionally, we evaluate on a version of this task that allows us to quantify repetition, i.e., the grammatical case of the elicited pronoun is the same as the case shown in the context. Examining model performance where a pronoun is shown in one grammatical case and then elicited in a different one would be interesting to probe syntactic generalization.

### v3

@@ Abstract
Robust, faithful and harm-free pronoun use for individuals is an important goal for language model development as their use increases, but prior work tends to study only one or two of these characteristics at a time. To measure progress towards the combined goal, we introduce the task of pronoun fidelity: given a context introducing a co-referring entity and pronoun, the task is to reuse the correct pronoun later. We present RUFF, a carefully-designed dataset of over 5 million instances to measure robust pronoun fidelity in English, and we evaluate 37 model variants from nine popular families, across architectures (encoder-only, decoder-only and encoder-decoder) and scales (11M-70B parameters). When an individual is introduced with a pronoun, models can mostly faithfully reuse this pronoun in the next sentence, but they are significantly worse with she/her/her, singular they and neopronouns. Moreover, models are easily distracted by non-adversarial sentences discussing other people; even one sentence with a distractor pronoun causes accuracy to drop on average by 34 percentage points. Our results show that pronoun fidelity is not robust, in a simple, naturalistic setting where humans achieve nearly 100% accuracy. We encourage researchers to bridge the gaps we find and to carefully evaluate reasoning in settings where superficial repetition might inflate perceptions of model performance.
@@ 1 Introduction
Accordingly, it is important for large language models (LLMs) to use pronouns faithfully and without causing harm. To this end, many studies have explored how LLMs handle pronouns, showing that they stereotypically associate pronouns and occupations (Kurita et al., 2019), reason about co-referring pronouns and entities better when they conform to stereotypes (Tal et al., 2022), fail when exposed to novel pronoun phenomena such as neopronouns (Lauscher et al., 2023), and cannot consistently reuse neopronouns during generation (Ovalle et al., 2023). These shortcomings create differences in quality of service and cause representational harm, amplifying discrimination against certain pronoun users (Blodgett et al., 2020; Dev et al., 2021).
In work on LLM pronoun use, a question that has gone unexamined thus far is: How robust is model faithfulness to pronouns when discussing more than one person? To answer this question, we propose pronoun fidelity2), a new task to investigate realistic model reasoning about pronouns, and we introduce RUFF3), a novel, large-scale dataset of over 5 million instances, to evaluate this task. With this dataset, we present an analysis of pronoun fidelity across 37 variants from nine popular language model families covering architectures and scales, to investigate whether models are reasoning, repeating, or just biased.
@@ 3 RUFF Dataset
Figure 2: Template assembly for RUFF: occupation-specific task templates are matched with generic context templates (introductions and optional distractors) that are instantiated with disjoint pronoun sets. This creates realistic but controlled narratives that allow us to measure robust pronoun fidelity.
To evaluate Robust pronoUn Fidelity at scale, we create RUFF, an evaluation dataset of narratives. Each dataset instance describes a simple narrative with 1-2 people, but rather than narrative data that focuses on commonsense event reasoning (Mostafazadeh et al., 2016), we focus on pronominal reasoning, like Rudinger et al. (2018).
Specifically, we examine four third-person pronouns in three grammatical cases (nominative, accusative and possessive dependent); in addition to the English masculine (he/him/his) and feminine (she/her/her) pronouns, we heed Lauscher et al.’s (2022) call for more inclusive NLP research by examining two more pronoun sets that are less well-studied in NLP: singular they (they/them/their), the pronoun of choice of over 75% of respondents to the Gender Census (Lodge, 2023), and xe/xem/xyr, the most popular neopronoun according to the same census. Our narratives cover 60 occupations and corresponding participants (see Appendix A), following Winogender schemas (Rudinger et al., 2018), as their bias characteristics are well-studied in NLP. In total, RUFF contains over 5 million data instances. Each instance is designed to have an unambiguous answer, and is constructed with a 3-step pipeline: template creation (§3.1), template assembly (§3.2), and data validation (§3.3).
@@ 3.1 Template creation
Below, we describe how we create occupation-specific task templates and generic context templates for introductions and distractors.
Task templates. We create one task sentence template per occupation and grammatical case, with an unambiguous, unique coreference between the pronoun and occupation, for a total of 180 templates. For instance, charges for preparing tax returns can only belong to an accountant, never a taxpayer, the corresponding participant.
Context templates. The ideal context template would be: (1) flexible across different occupations and participants, for a controlled setting to test robustness; (2) cohesive in a multi-sentence narrative leading up to the task template about an occupation; and (3) neutral, not dramatically affecting the prediction of a certain pronoun. Templates such as He is an accountant are well-established for testing word embedding associations (Caliskan et al., 2017; May et al., 2019). They are flexible and neutral (they are even referred to as “semantically bleached” templates in the literature), but it is unnatural to use more than one consecutively. Natural corpora like Levy et al. (2021) have the most potential for creating cohesive narratives, but contain occupation-specific sentences that are inflexible and sometimes also non-neutral, e.g., ungrammatical with singular they.
For a setting that satisfies all three criteria, we create context templates with generic themes, e.g., universal human emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.). The generic themes make them flexible for use across all occupations and participants. Templates of the same polarity can be stacked into a cohesive narrative, e.g., a narrative about a taxpayer having a bad day after sleeping poorly and missing a meal. Our templates are created to be grammatical with all pronoun sets we consider, which satisfies neutrality. Additionally, our use of both positive and negative versions of templates (i.e., happy and unhapppy) as well as our variety of templates allows us to mitigate potential implicit biases when aggregated (Alnegheimish et al., 2022).
To reflect natural and coherent use of pronouns in discourse, we create explicit (definite reference + pronoun) and implicit (pronoun-only) versions of 10 context templates per grammatical case, for a total of 30 templates. Each explicit context template begins with an entity and introduces the pronoun in a clause, e.g., The taxpayer needed coffee because their day had started very early, while implicit templates are simple sentences like Their sleep had been fitful. See Appendix B for more detail on template creation and assembly.
@@ 3.2 Template assembly
Figure 2 shows how we instantiate and combine templates to assemble our data instances: first, we select an occupation ($e_{a}$) and one of its task templates. We pick a pronoun ($p_{a}$) to use as ground truth and instantiate a random context template with the selected occupation and pronoun. The simplest version of the pronoun fidelity task includes just this introduction sentence followed by the task sentence. Instantiating 10 templates with 4 different pronoun sets and pairing them with task templates for 60 occupations across 3 grammatical cases gives us a total of 7,200 unique instances for this version of the task.
To create more complex data instances, we insert a variable number of distractor sentences between the introduction and task sentences, discussing a participant $e_{b}$ with a different pronoun $p_{b}$. These are also sampled from the set of context templates (see Appendix B for details). Instantiating 4 templates with 3 previously unused pronouns gives 86,400 unique instances with one distractor.
@@
FLAN-T5 small (77M), base (248M), large (783M), xl (2.85B), xxl (11.3B) Encoder-decoder
@@
Table 2: Models we experiment with across a range of sizes (11M-70B parameters) and architectures.
@@ 3.3 Data validation
We validate all task and context templates. To verify that the pronoun fidelity task is easy and unambiguous for humans, and to create a ceiling for model evaluation, we also validate a subset of task instances with 0-5 distractors. Annotator information is shown in Appendix C and all annotator instructions are provided in Appendix D.
@@ 4 Experimental Setup
Figure 3: Model evaluation overview: pseudo log likelihoods (PLLs) and log likelihoods (LLs) of verbalized instances are used for encoder-only and decoder-only models; generations are used for chat models.
We list our models, evaluation methods, and metrics. Further details are provided in Appendix E.
@@ 4.1 Models
We experiment with 37 transformer-based language model variants from nine popular model families (see Table 2), which we chose to evaluate the effects of architecture and scaling. Our encoder-only models are from the BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), ALBERT-v2 (Lan et al., 2020) and MosaicBERT (Portes et al., 2023) model families, as the first three remain well-used in NLP, and the last is trained on much more data. As for our decoder-only models, we select the popular Llama-2 (Touvron et al., 2023) model family, as well as OPT (Zhang et al., 2022) and Pythia (Biderman et al., 2023) for their large range of model sizes. In Appendix H, we also experiment with popular chat models that are further trained with instruction-tuning and reinforcement learning, to evaluate task performance with prompting; specifically, we use decoder-only Llama-2-chat models (Touvron et al., 2023) and encoder-decoder FLAN-T5 models (Chung et al., 2024).
@@ 4.2 Obtaining predictions
Figure 3 shows an overview of our evaluation methods. Decoder-only and encoder-only models are evaluated comparably in a forced choice setting: following Hu and Levy (2023), we take direct measurements of probabilities as a proxy for models’ metalinguistic judgements. Generations are obtained from chat models and post-processed to obtain unique pronouns, if any.
Encoder-only and decoder-only models. We verbalize four versions of each data instance, i.e., we fill in the blank with each of the four pronouns we consider, creating four options. We then obtain model probabilities for each of these four options, and select the highest probability option as the model’s choice. We use log likelihoods for decoder-only models and pseudo log likelihoods for encoder-only models, following prior work (Salazar et al., 2020; Kauf and Ivanova, 2023). We do not use masked token prediction due to tokenization issues with neopronouns (Ovalle et al., 2024); briefly, we want xe to be tokenized “normally” (which is often as two tokens) rather than a single UNK token.
Chat models. Following common practice, we evaluate chat models (FLAN-T5 and Llama-2-chat) using vanilla and chain-of-thought prompting. Following Sclar et al. (2024), we show the range of expected performance with 10 different prompts, inspired by the prompts to elicit coreferences in the FLAN collection (Longpre et al., 2023). See Appendix F for more methodological details and Appendix H for results.
@@ 5 Model Predictions with No Context
As we cannot evaluate accuracy on a task with no single correct answer, we show the counts of model predictions of different pronoun declensions in Figure 4, averaged over all models. Model-specific counts are shown in Appendix G. Even though our task sentences are designed such that any pronoun set can be used grammatically, all models tend to assign higher probability to he/him/his than other pronoun sets.
@@ 6 Injecting an Introductory Context
Which pronouns are harder? Even in the simplest case of the pronoun fidelity task, patterns emerge when split by pronoun, as shown in Figure 6. Overall model accuracy on he/him/his is significantly higher than she/her/her, which in turn is significantly higher than both they/them/their and xe/xem/xyr, in line with previous findings that language technology has gaps when it comes to neopronouns (Lauscher et al., 2023). Models show intriguing patterns with these last two pronoun sets. Most encoder-only models appear to handle the neopronoun better than singular they (e.g., BERT-large has an accuracy of 0.78 on xe/xem/xyr compared to 0.60 on they/them/their), which warrants further investigation. Decoder-only models smaller than 6.7B parameters struggle with the neopronoun, with every OPT and Pythia model smaller than 2.7B parameters performing below chance, and in some cases (e.g., Pythia-14M, Pythia-70M and Pythia-160M) even performing close to 0.0. Beyond this scale, however, models perform better on xe/xem/xyr than on singular they, with Llama-13B achieving 0.96 accuracy on the neopronoun. These differences are statistically significant. As the training data for individual model families is the same, this might suggest that decoder-only models generalize to novel pronouns starting at the scale of 6.7B parameters, but in light of Schaeffer et al. (2023), this result could just as well be a mirage resulting from our use of accuracy, a discontinuous metric. In either case, our observations could also explain the poor performance that some previous studies of neopronouns find, as the largest model that Hossain et al. (2023) experiment with, for instance, is OPT-6.7B. The lower performance of bigger models with singular they could also be a reflection of human processing difficulties with definite, specific singular they, as has been observed in linguistics (Conrod, 2019).
@@ 7 Adding Distractors
Using vanilla and chain-of-thought prompting (Appendix H) show the same patterns of degradation, reinforcing that model pronoun fidelity is not robust, and good performance with no distractors (§6) is likely not due to “reasoning” at all.
@@ 8 Distractibility versus Bias
In adding distractor sentences, we add distance from the introduction via additional tokens that might make the model forget the original occupation-pronoun association, and the distractor pronoun also acts as a competing token that the model might accidentally repeat. In this section, we focus on the error cases to disentangle whether models are “forgetting” and reverting to biased predictions from Section 5, or if they are actually being distracted. When a model gets the answer wrong, it is for one of three reasons: (1) distractibility, i.e., repeating the distractor pronoun, (2) bias, i.e., reverting to the model’s context-free prediction, or (3) picking a completely different pronoun. Our example illustrates all three possibilities, and we hypothesize that the first two possibilities are much more frequent than the third.
@@
We first examine model distractibility, i.e., what percentage of errors are caused by models repeating the distractor pronoun instead of the correct pronoun. As expected, Figure 9 shows that across models, distraction is indeed the primary type of error for most models. Decoder-only models get increasingly distracted with more distractors, i.e., the proportion of errors due to distractor pronoun repetition steadily increases as distractors are added, saturating just below 85%. On the other hand, encoder-only models seem to become less distractible with the addition of more distractors. We know from the previous section that encoder-only models recover in their pronoun fidelity with 2-5 distractors, but here we measure distractibility as a percentage of all errors. Thus, a constant or increasing proportion of all the model errors could be due to distraction, and the fact that it isn’t for encoder-only models is quite surprising! We leave it to future work to investigate whether this behaviour relates to positional bias or context use.
@@
The high distractibility of all models shows that models are not robust reasoners, and the contrast in error behaviour between encoder-only and decoder-only models further highlights their differences. This shows that claims about decoder-only models should not be applied to all LLMs, and that reasoning must be evaluated carefully, accounting for the possibility of inflated performance due to shallow heuristics like repetition.
@@ 9 Discussion and Future Work
Our results show that even the biggest models of today are not up to the task of pronoun fidelity once it includes a single sentence discussing another person. All models are easily distracted, but encoder-only models and decoder-only models show very different patterns both in performance degradation with more distractors and their reasons for errors. Performance on this type of reasoning task should be evaluated carefully, with attention to how the overall patterns break down by different pronouns, and accounting for the possibility of repetition. Below we expand on some questions raised by our findings.
@@
Improving robust pronoun fidelity. A natural direction of future work is to solve the problem of robust pronoun fidelity, particularly in decoder-only models, which are unlikely to be replaced by encoder-only models with poorer generation abilities. A promising direction might be to encourage models to explicitly track associations between individuals and pronoun sets, just as people do. In fact, prior work has noted success with generative models when explicitly tracking mentions of entities across multiple tasks (Ji et al., 2017) and in the context of story generation (Fan et al., 2019). We urge researchers interested in this direction to treat RUFF as an evaluation dataset, as it was designed. Due to the presence of positional and associative heuristics (see Appendix 12), RUFF should not be seen as a source of data for fine-tuning or in-context learning, which is also why we do not run these experiments.
@@
On “reasoning.” Throughout the paper, we refer to “reasoning,” but this is inaccurate. Even the higher performance of encoder-only models cannot accurately be attributed to “reasoning” in the same way that we use this word for humans, as these models are not grounded in meaning from the real world (Bender and Koller, 2020). We use the word reasoning in line with other work in the field, but note that as these are all language models, it is more accurate to say that the way that decoder-only models model language is prone to repetition—or stochastic parroting (Bender et al., 2021)—of recent examples of the same word class, compared to encoder-only models.
@@
Why exactly do we see the patterns we see? Our dataset design and error analysis shed light on model behaviour, allowing us to evaluate different architectures comparably and disentangle the effects of repetition, distraction and statistical bias. However, it is beyond the scope of this paper to investigate where in the model architecture, neurons or pre-training data this comes from and what we can do about it towards improving reasoning and mitigating bias. Tools for model interpretability, e.g., attribution analysis, could help here, and are an important direction for future work.
@@
Beyond our dataset. Given the breadth of our task definition, future work could examine pronoun fidelity in other contexts, e.g., for participants, for names by extending Hossain et al. (2023), with differently ordered sentences, with real-world data as in Webster et al. (2018) and Levy et al. (2021), and in domains beyond simple narratives (Pradhan et al., 2013). Additionally, we evaluate on a version of this task that allows us to quantify repetition, i.e., the grammatical case of the elicited pronoun is the same as the case shown in the context. Examining model performance where a pronoun is shown in one grammatical case and then elicited in a different one would be interesting to probe syntactic generalization.